OpenStudy (migitmack):

the sum of the length of the two bases of a trapezoid is 220cm and its area is 66 cm. what is the height of the trapezoid. I know I know how to do this but I cant seem to get it.

OpenStudy (hexagon001):

0.5(a+b)h=area of trapeziod. a+b=220 and area =66 so we have.. 0.5(220)h=66 110h=66 h=66/110=3/5=0.6

OpenStudy (migitmack):

oh wow XD thank you